OpenStudy (anonymous):

Does anyone how to convert Celsius into Fahrenheit? such as 98.6F into Celsius? thanks!

OpenStudy (anonymous):

or what formula its used?

OpenStudy (anonymous):

The formulas are freely available on wikipedia: http://en.wikipedia.org/wiki/Celsius

OpenStudy (anonymous):

sorry I don't Trust Wikpedia..

OpenStudy (anonymous):

Then you shouldn't trust me either. Or anyone on this site. We're the Illuminati here.

OpenStudy (amistre64):

you can reconstruct the equations by using 2 points

OpenStudy (amistre64):

(F,C): (32,0) and (212,100) F and C form a linear relationship such that their conversions form a line

OpenStudy (amistre64):

(32,0) and (212,100) -32 -32 ------ -------- 0,0 180,100 slope of the line is 10/18 = 5/9 C = 5/9 (F - 32)

OpenStudy (anonymous):

okay amistre thanks!

OpenStudy (anonymous):

purple I needed someone to show me, an example, not just give me a website... but thnks for your help anyway...

OpenStudy (amistre64):

hope it helps ;)